# AdaptEval: Evaluating Large Language Models on Domain Adaptation for Text Summarization

### v2

@@ Abstract
Despite the advances in the abstractive summarization task using Large Language Models (LLM), there is a lack of research that asses their abilities to easily adapt to different domains. We evaluate the domain adaptation abilities of a wide range of LLMs on the summarization task across various domains in both fine-tuning and in-context learning settings. We also present AdaptEval, the first domain adaptation evaluation suite. AdaptEval includes a domain benchmark and a set of metrics to facilitate the analysis of domain adaptation. Our results demonstrate that LLMs exhibit comparable performance in the in-context learning setting, regardless of their parameter scale.
@@ 1 Introduction
In this paper, we assess the domain adaptation abilities of 11 models, including conventional encoder-decoder models and a wide range of LLMs in various parameter sizes, on the summarization task. In particular, we experiment with the fine-tuning and in-context learning (ICL) settings and evaluate their performance across various domains (i.e. governmental, medical, and scientific), reporting scores on a collection of automatic—ROUGE Lin (2004) and BERTScore Zhang et al. (2019)—and domain adaptation metrics. The latter includes domain vocabulary overlap Yu et al. (2021), and our adaptations of G-eval Liu et al. (2023) and token distribution shift Lin et al. (2023) to the task.
@@ 2.2 Evaluation Metrics
G-eval uses GPT-4 OpenAI (2023) with chain-of-thought prompting Wei et al. (2022) to evaluate summaries across quality features, such as coherence or fluency, achieving high correlation with human judgments Liu et al. (2023). Similarly, we design a prompt to score the degree to which a summary adheres to the domain language on a scale from 1 to 5. Our prompt includes the reasoning steps generated by GTP-4 as in Liu et al. (2023) (see Appendix D).
@@
ChatGPT 0.841 38.58 22.92 0.837 38.39 23.15 x x x
Llama2  7b 0.819 35.95 21.11 0.824 35.34 20.92 x x x
Llama2 70b 0.845 37.61 22.40 0.842 36.65 23.03 x x x
Mistral 7b 0.816 32.05 21.30 0.802 23.61 17.76 x x x
Vicuna  7b 0.831 36.29 21.54 0.827 34.65 20.31 x x x
@@
Table 3: BERTScore F1, DVO (%), and the geometric mean of ROUGE-1/2/L (ROUGE) of all models across the three domains. The value ‘x’ implies that the model was not evaluated under those settings. 1/2/3 indicate fine-tuning with 1k, 5k, and 10k instances, respectively. See complete table in Appendix B.
@@ 3 Domain Adaptation Task
We assess the performance of 11 models across domains in both fine-tuning and ICL settings. [^4] We exclude GovReport from the in-context learning setting with LLMs, since the documents are too long to fit into the model’s context window.
@@ 3.1 Models Selection
Science Medical
base 2-shot KL TSR KL TSR
Llama2  7b vs. 7b 19.70 92.14 19.27 97.44
Llama2 13b vs. 13b 15.58 96.95 16.53 96.76
Llama2 70b vs. 70b 16.78 95.68 17.12 98.19
Vicuna  7b vs. 7b 17.67 92.35 18.32 93.89
Vicuna 13b vs. 13b 18.12 97.13 17.34 90.70
Mistral 7b vs. 7b 13.88 91.33 14.01 95.40
2-shot 2-shot KL TSR KL TSR
Llama2 70b vs. 7b 0.43 3.92 0.46 5.01
Llama2 13b vs. 7b 0.21 2.87 0.38 1.67
Llama2 70b vs. 13b 0.47 5.18 0.31 3.50
Vicuna 13b vs. 7b 0.25 2.07 0.38 4.57
base FT KL TSR KL TSR
Llama2  7b vs. 7b 0.81 12.40 0.35 4.70
Llama2 13b vs. 13b 0.51 6.84 0.48 7.32
Mistral 7b vs. 7b 0.52 11.54 0.37 4.42
Table 4: Effect of different model sizes, two-shot in-context learning, and Fine-Tuning in terms of token distribution shift scores—KL divergence and Token Shift Rate (%) calculated over 10 samples. Two-shot has a major impact on the model’s predictions. The low scores between different model sizes indicate that parameter size does not have a significant effect on domain adaptation in the two-shot setting. (See complete table in Appendix C.
@@ 3.2 Results
To confirm these findings, we also compute G-eval scores, which have a strong correlation with human judgments, along with our addition to measure domain adaptation, on a random sample of 25 articles from arXiv and PubMed. [^5] Due to the costs of using GPT-4 with large prompts, we only report the scores of three quality features (coherence, fluency, and domain adaptation) on four models outputs of 25 random instances from arXiv and PubMed. We do not report scores on GovReport since we did not evaluate it under the two-shot setting with LLMs. The scores on arXiv data are consistent with our previous results, showing that ICL achieves the best performance, and the model parameter size does not have a significant impact. However, PubMed obtains remarkably low scores, which highlights the difficulty of the models to adapt to the medical domain.
@@
2-shot arXiv PMed arXiv PMed arXiv PMed
Llama2  7b 4.20 1.0 3.80 2.0 2.72 2.0
Llama2 70b 3.96 1.0 3.20 1.0 2.56 1.0
@@
Llama2  7b 3.48 2.0 2.08 2.0 2.04 2.0
PEGASUS-X 3.88 2.8 2.88 2.0 2.40 2.0
Table 5: G-Eval scores on 25 random samples from the arXiv and PubMed datasets in terms of coherence (1-5), fluency (1-3), and our Domain Adaptation (DA) (1-5).
@@ 3.3 Manual Evaluation
Two in-house domain experts perform a blind manual evaluation of the same arXiv sample used in G-eval evaluation (Table 5). Specifically, the sample comprises 25 random arXiv articles paired with four different summaries generated with Llama2 (7b and 70b) in the two-shot setting, fine-tuned Llama2 (7b) and PEGASUS-X. To avoid biases, we randomly shuffle the evaluation instances and their summaries for each annotator.
@@ 5 Conclusion
We evaluate the domain adaptation abilities of Large Language Models across scientific, medical, and governmental domains using a set of adapted evaluation metrics. Additionally, we release AdaptEval, an evaluation suite that facilitates the analysis of domain adaptation. Our experiments show that smaller LLMs exhibit domain-shift challenges, but they are able to achieve comparable performance to larger LLMs when provided with only two learning examples. In contrast, fine-tuning does not have a significant impact on the vocabulary domain, but only on stylistic tokens. Overall, the G-eval scores indicate that the medical domain is challenging for these models. We expect our work to encourage and facilitate further research on domain adaptation with LLMs across domains.
@@ A.1 Zero-shot Setting
We used the instruct-tuned or chat versions of the models. As for ChatGPT, we used the OpenAI API [^6] https://platform.openai.com/ and the latest snapshot available, gpt-3.5-turbo-0613 from June 13th, 2023. For zero-shot setting, we used Llama2 (7b) [^7] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf/, Llama2 (13b) [^8] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf/, Llama2 (70b)  [^9] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf/, Vicuna (7b) [^10] https://huggingface.co/lmsys/vicuna-7b-v1.5, Vicuna (13b) [^11] https://huggingface.co/lmsys/vicuna-13b-v1.5, Falcon (7b) [^12] https://huggingface.co/tiiuae/falcon-7b, Falcon (40b) [^13] https://huggingface.co/tiiuae/falcon-40b, and Mistral AI (7b) [^14] https://mistralai/Mistral-7B-Instruct-v0.1. When generating summaries, we sample a maximum of 256 tokens for the arXiv and PubMed datasets, while scaling to 1024 tokens for the GovReport dataset, as is standard procedure in other contemporary publications. The prompts used 0-shot and 2-shot settings for generating the summaries is shown in Table 9.
@@ A.3 Fine-tuning Setting
Appendix B Automatic Scores and DVO
Table 7 shows the performance of all models across domains in terms of ROUGE, BertScores, and Domain Vocabulary Overlap (DVO).
Medical Science Government
BERTScore DVO ROUGE BERTScore DVO ROUGE BERTScore DVO ROUGE
Zero-shot Setting
ChatGPT 0.844 36.69 24.81 0.838 36.58 23.95 0.859 37.73 22.34
PEGASUS-X 0.690 6.28 3.55 0.538 11.98 5.85 0.736 5.58 9.06
Falcon  7b 0.811 31.87 13.68 0.810 30.16 14.54 0.821 31.49 13.86
Falcon 40b 0.816 35.51 13.85 0.822 34.98 17.59 0.827 35.51 13.85
Llama2  7b 0.783 21.15 10.94 0.818 28.61 18.33 0.845 34.36 18.86
Llama2 13b 0.764 20.78 6.26 0.783 23.48 8.58 0.797 24.04 10.80
Llama2 70b 0.842 35.50 24.59 0.837 35.22 23.35 0.855 36.05 21.48
Mistral 7b 0.788 24.78 9.44 0.806 28.81 13.68 0.815 31.18 12.02
Vicuna  7b 0.727 9.49 2.11 0.781 23.94 7.93 0.813 30.69 10.80
Vicuna 13b 0.745 15.76 1.58 0.763 19.07 4.43 0.783 27.18 7.17
Two-shot Setting
ChatGPT 0.841 38.58 22.92 0.837 38.39 23.15 x x x
Llama2  7b 0.819 35.95 21.11 0.824 35.34 20.92 x x x
Llama2 13b 0.820 35.02 19.00 0.809 32.30 18.97 x x x
Llama2 70b 0.845 37.61 22.40 0.842 36.65 23.03 x x x
Mistral 7b 0.816 32.05 21.30 0.802 23.61 17.76 x x x
Vicuna  7b 0.831 36.29 21.54 0.827 34.65 20.31 x x x
Vicuna 13b 0.822 35.51 19.69 0.807 33.32 14.86 x x x
Fine-tuning Setting
BART 0.852 37.03 24.80 0.844 34.15 22.20 0.856 25.14 28.44
PEGASUS-X 0.850 28.72 31.18 0.852 34.61 28.11 0.868 22.07 31.98
Llama2  7b1 0.859 33.61 25.81 0.858 33.06 25.30 0.850 29.30 24.81
Llama2  7b2 0.861 35.15 26.00 0.856 30.49 25.46 x x x
Llama2  7b3 0.862 33.71 26.81 0.854 27.43 25.35 x x x
Llama2 13b2 0.862 35.28 26.26 0.860 32.67 26.47 x x x
Mistral 7b2 0.863 35.81 27.17 0.863 34.00 27.29 0.833 21.66 23.08
Appendix C Token Distribution Shift
Science Medical Government
base 2-shot KL TSR KL TSR KL TSR
Llama2  7b vs. 7b 19.70 92.14 19.27 97.44 x x
Llama2 13b vs. 13b 15.58 96.95 16.53 96.76 x x
Llama2 70b vs. 70b 16.78 95.68 17.12 98.19 x x
Vicuna  7b vs. 7b 17.67 92.35 18.32 93.89 x x
Vicuna 13b vs. 13b 18.12 97.13 17.34 90.70 x x
Mistral 7b vs. 7b 13.88 91.33 14.01 95.40 x x
2-shot 2-shot KL TSR KL TSR KL TSR
Llama2 70b vs. 7b 0.43 3.92 0.46 5.01 x x
Llama2 13b vs. 7b 0.21 2.87 0.38 1.67 x x
Llama2 70b vs. 13b 0.47 5.18 0.31 3.50 x x
Vicuna 13b vs. 7b 0.25 2.07 0.38 4.57 x x
base FT KL TSR KL TSR KL TSR
Llama2  7b vs. 7b 0.81 12.40 0.35 4.70 21.49 15.15
Llama2 13b vs. 13b 0.51 6.84 0.48 7.32 x x
Mistral 7b vs. 7b 0.52 11.54 0.37 4.42 0.18 3.21
Table 8: Effect of different model sizes, two-shot in-context learning, and Fine-Tuning in terms of token distribution shift scores—KL divergence and Token Shift Rate (%) calculated over 10 samples. Two-shot has the major impact on the models’ predictions. The low scores between different model sizes indicate that parameter size does not have a significant effect on domain adaptation in the two-shot setting. (See complete table in Appendix C.
Table 8 demonstrates the Token Distribution Shift between the base and our Two-shot prompt and Fine-tuned models.
Appendix D LLM Prompting
Table 9 and Table 10 illustrate the prompts used to generate summaries and to score the domain adaptation of summaries using GPT-4, respectively.

### v3

@@ Abstract
Despite the advances in the abstractive summarization task using Large Language Models (LLM), there is a lack of research that assess their abilities to easily adapt to different domains. We evaluate the domain adaptation abilities of a wide range of LLMs on the summarization task across various domains in both fine-tuning and in-context learning settings. We also present AdaptEval, the first domain adaptation evaluation suite. AdaptEval includes a domain benchmark and a set of metrics to facilitate the analysis of domain adaptation. Our results demonstrate that LLMs exhibit comparable performance in the in-context learning setting, regardless of their parameter scale.
@@ 1 Introduction
In this paper, we assess the domain adaptation abilities of 11 models, including conventional encoder-decoder models and a wide range of LLMs in various parameter sizes, on the summarization task. In particular, we experiment with fine-tuning and in-context learning (ICL) settings and evaluate their performance across various domains (i.e. governmental, medical, and scientific), reporting scores on a collection of automatic—ROUGE Lin (2004) and BERTScore Zhang et al. (2019)—and domain adaptation metrics. The latter includes domain vocabulary overlap Yu et al. (2021), and our adaptations of G-eval Liu et al. (2023) and token distribution shift Lin et al. (2023) to the task.
@@ 2.2 Evaluation Metrics
G-eval uses GPT-4 OpenAI (2023) with chain-of-thought prompting Wei et al. (2022) to evaluate summaries across quality features, such as coherence or fluency, achieving high correlation with human judgments Liu et al. (2023). Similarly, we design a prompt to score the degree to which a summary adheres to the domain language on a scale from 1 to 5. Our prompt includes the reasoning steps generated by GTP-4 as in Liu et al. (2023) (see Appendix B).
@@
Llama2  7b 0.819 35.95 21.11 0.824 35.34 20.92 0.847 30.22 17.39
Mistral 7b 0.816 32.05 21.30 0.802 23.61 17.76 0.844 30.08 19.21
Vicuna  7b 0.831 36.29 21.54 0.827 34.65 20.31 0.851 30.28 17.29
Llama2 13b 0.820 35.02 19.00 0.809 32.30 18.97 0.814 29.92 14.30
Vicuna 13b 0.822 35.51 19.69 0.807 33.32 14.86 0.789 29.34 8.34
Llama2 70b 0.845 37.61 22.40 0.842 36.65 23.03 0.851 29.59 18.72
ChatGPT 0.841 38.58 22.92 0.837 38.39 23.15 0.853 30.44 16.82
GPT-4o mini 0.842 30.64 23.18 0.835 29.14 21.47 0.850 30.40 16.04
@@
Llama2 13b2 0.862 35.28 26.26 0.860 32.67 26.47 x x x
@@ 3 Domain Adaptation Task
We assess the performance of 11 models across domains in both fine-tuning [^4] We exclude GovReport from fine-tuning on 5k and 10k samples, since the train set doesn’t have enough documents to fit into the models context window of 4096 tokens—only 1148 instances with maximum 4k length in the training split. and ICL settings.
@@ 3.1 Models Selection
Science Medical Government
base 2-shot KL TSR KL TSR KL TSR
Llama2  7b vs. 7b 19.70 92.14 19.27 97.44 17.40 94.33
Mistral 7b vs. 7b 13.88 91.33 14.01 95.40 13.40 90.00
Vicuna  7b vs. 7b 17.67 92.35 18.32 93.89 15.42 94.04
Llama2 13b vs. 13b 15.58 96.95 16.53 96.76 14.67 98.82
Vicuna 13b vs. 13b 18.12 97.13 17.34 90.70 16.79 99.10
Llama2 70b vs. 70b 16.78 95.68 17.12 98.19 13.10 92.36
2-shot 2-shot KL TSR KL TSR KL TSR
Llama2 13b vs. 7b 0.21 2.87 0.38 1.67 0.32 10.38
Vicuna 13b vs. 7b 0.25 2.07 0.38 4.57 0.24 0.00
Llama2 70b vs. 13b 0.47 5.18 0.31 3.50 0.49 4.92
Llama2 70b vs. 7b 0.43 3.92 0.46 5.01 0.54 6.88
base FT KL TSR KL TSR KL TSR
Llama2  7b vs. 7b 0.81 12.40 0.35 4.70 21.49 15.15
Mistral 7b vs. 7b 0.52 11.54 0.37 4.42 0.18 3.21
Llama2 13b vs. 13b 0.51 6.84 0.48 7.32 x x
Table 4: Effect of different model sizes, two-shot in-context learning, and Fine-Tuning in terms of token distribution shift scores—KL divergence and Token Shift Rate (%) calculated over 10 samples. Two-shot has the major impact on the models’ predictions. The low scores between different model sizes indicate that parameter size does not have a significant effect on domain adaptation in the two-shot setting.
@@ 3.2 Results
To confirm these findings, we also evaluate the summaries using GPT-4 shown in Table 5, which have a strong correlation with human judgments, along with our addition to measure domain adaptation, on a random sample of 25 articles. [^5] Due to the costs of using GPT-4 with large prompts, we only report the scores on four models outputs of 25 random instances. The scores on arXiv data are consistent with our previous results, showing that ICL achieves the best performance, and the model parameter size does not have a significant impact. However, PubMed obtains remarkably low scores, which highlights the difficulty of the models to adapt to the medical domain. The LLMs however, find it easier to adapt to the Government domain.
@@
2-shot arXiv PubMed GovReport arXiv PubMed GovReport arXiv PubMed GovReport
Llama2  7b 4.20 1.0 4.04 3.80 2.0 3.96 2.72 2.0 2.96
Llama2 70b 3.96 1.0 4.40 3.20 1.0 3.96 2.56 1.0 3.00
@@
Llama2  7b 3.48 2.0 4.16 2.08 2.0 3.40 2.04 2.0 2.84
PEGASUS-X 3.88 2.8 4.40 2.88 2.0 3.72 2.40 2.0 2.72
Table 5: Evaluation scores using GPT-4 on 25 random samples from the arXiv, PubMed and GovReport datasets in terms of coherence (1-5), fluency (1-3), and our Domain Adaptation (DA) (1-5).
@@ 3.3 Manual Evaluation
Two in-house domain experts perform a blind manual evaluation of the same arXiv samples used in GPT-4 evaluation (Table 5). The setting comprises of 25 random arXiv articles paired with four different summaries generated with Llama2 (7b and 70b) in the two-shot setting, fine-tuned Llama2 (7b) and PEGASUS-X. To avoid biases, we randomly shuffle the evaluation instances and their summaries for each annotator.
@@ 5 Conclusion
We evaluate the domain adaptation abilities of Large Language Models across scientific, medical, and governmental domains using a set of adapted evaluation metrics. Additionally, we release AdaptEval, an evaluation suite that facilitates the analysis of domain adaptation. Our experiments show that smaller LLMs exhibit domain-shift challenges, but they are able to achieve comparable performance to larger LLMs when provided with only two learning examples. In contrast, fine-tuning does not have a significant impact on the vocabulary domain, but only on stylistic tokens. Overall, the G-eval scores indicate that the medical domain is challenging for these models. We expect our work to encourage and facilitate further research on domain adaptation with LLMs across domains. We plan to continue this research in future work.
Limitations
To fairly compare the performance of the different models, we generally restricted our evaluation to those models with context window of 4096. An exception is the language model BART with a context window of 1024. Additionally, due to the high costs of performing human evaluations on multiple domains, we only annotated ArXiv data to reaffirm the results obtained through the automatic metrics. Our goal is to facilitate the evaluation of models across domains to the research community. Therefore, our suite consists of a set of metrics to evaluate domain adaptation and general summarization quality, allowing for a comprehensive comparison of the models performance on multiple datasets. Lastly, given the cost associated with GPT-4, we performed LLM-based evaluation on only 25 random samples.
@@
Acknowledgements
This project is supported by Ringier, TX Group, NZZ, SRG, VSM, viscom, the ETH Zurich Foundation, and the German Federal Ministry of Education and Research (BMBF) grant 01IS17049 Software Campus 2.0 (TU München).
@@ A.1 Zero-shot Setting
We used the instruct-tuned or chat versions of the models. As for ChatGPT, we used the OpenAI API [^6] https://platform.openai.com/ and the latest snapshot available, gpt-3.5-turbo-0613 from June 13th, 2023. For zero-shot setting, we used Llama2 (7b) [^7] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf/, Llama2 (13b) [^8] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf/, Llama2 (70b)  [^9] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf/, Vicuna (7b) [^10] https://huggingface.co/lmsys/vicuna-7b-v1.5, Vicuna (13b) [^11] https://huggingface.co/lmsys/vicuna-13b-v1.5, Falcon (7b) [^12] https://huggingface.co/tiiuae/falcon-7b, Falcon (40b) [^13] https://huggingface.co/tiiuae/falcon-40b, and Mistral AI (7b) [^14] https://mistralai/Mistral-7B-Instruct-v0.1.
When generating summaries, we sample a maximum of 256 tokens for the arXiv and PubMed datasets, while scaling to 1024 tokens for the GovReport dataset, as is standard procedure in other contemporary publications. The prompts used 0-shot and 2-shot settings for generating the summaries is shown in Table 7.
@@ A.3 Fine-tuning Setting
Appendix B LLM Prompting
Table 7 and Table 8 illustrate the prompts used to generate summaries and to score the domain adaptation of summaries using GPT-4, respectively. For evaluation, we use the prompts introduced by  Liu et al. (2023) for Coherence and Fluency. However, we craft our own prompt that asseses model’s ability to adapt to a new domain by evaluating the generated summaries.
@@ Appendix B LLM Prompting
Appendix C Sample Summaries
Table 9 shows the summaries generated by Llama2 7b under zero-shot, two-shot and fine-tuning setting.
Zero-shot Llama2 7b
mirnas are small non coding rnas that play important roles on the onset and progression of cancer . distinct mirnas regulate various cellular processes including cell cycle regulation , apoptosis , migration and invasion . altered expression of mirnas has been observed in breast cancer and is believed to play important roles including cell proliferation regulation , cell survival and tumorigenesis . furthermore , mirnas have been shown to play a critical role in cancer stem cell maintenance and tumorigenesis . emerging evidence suggests that mirnas may also be a attractive therapeutic target for cancer . understanding the specific functionalities of distinct mirnas and their role in disease may provide a new approach for cancer diagnosis and potentially new therapeutic strategies. reference listing: lee , j.a. (2011). mi rna : a duel role player in breast cancer progression . breast cancer : a global view . springer . 10.10079783642-135704_19 . 19 . 277 286 . abstract: micro rnas
Two-shot Llama2 7b
Small non-coding RNAs, such as siRNA, miRNA, piRNA, and snRNA, are transcribed into mRNA but remain untranslated in eukaryotic cells. They play a crucial role in regulating gene expression and have been implicated in the onset and progression of tumorigenesis. Emerging evidence suggests that small non-coding RNAs may represent a novel cancer therapy. Unlike mRNA, small non-coding RNAs are transcribed but never translated. Some small non-coding RNAs are transcribed from non-coding regions between genes, while others are transcribed together with coding mRNAs from the coding region of the genome. Altered expression of small non-coding RNAs has been observed in breast cancer, resulting in aberrant gene expression. Studies have shown that a high frequency ( 72.8%) of gene copy number abnormality in small non-coding RNA-containing regions occurs in human breast cancer. Decreased expression of certain small non-coding RNAs, such as mir-17/20, has been
Fine-tuned Llama2 7b
mirnas are small non - coding rnas that regulate gene expression and tumorigenesis . mirnas are involved in the regulation of the cell cycle , cellular proliferation , metastasis and drug resistance of breast cancer . the expression of mirnas is dysregulated in breast cancer . the dysregulated mirnas may serve as novel biomarkers for breast cancer . mirnas may serve as targets for gene therapy either alone or as an adjuvant to conventional therapy .
Table 9: Summaries generated by Llama2 7b under zero-shot, two-shot and fine-tuning setting for a sample article (id = 2) from PubMed test set.